1/1

Lord George Gordon Byron ,  Lord George Gordon Byron ,
She Walks in Beauty  Hon går i skönhet

tradukita de Theodor Lind

She walks in Beauty, like the Night Of cloudless climes and starry skies; And all that’s best of dark and bright Meet in her aspect and her eyes; Thus mellowed to the tender light Which Heaven to gaudy day denies – Hon går i skönhet, lik en natt, Med himmel klar och stjerners glans; Af ljus och skugga mötas gladt I min och blick, hvad skärast fanns, Så vekt, så skönt, så mildradt, att Dess glans af dagrar blott ej vanns.

One shade the more – one ray the less Had half impaired the nameless grace Which waves in every raven tress Or softly lightens o’er her face Where thoughts serenely sweet express How pure – how dear their dwelling place!

En stråle bort, en skugga mer, Och borta vore det behag, Som i den mörka locken ler, Som blickar ur de ljufva drag, Der hvarje tanke klart man ser, Hur ren den går i öppen dag.

And on that cheek, and o’er that brow, So soft – so calm – yet eloquent, The smiles that win, the tints that glow, But tell of days in goodness spent – A mind at peace with all below – A Heart – whose love is innocent!

Och panna, ögon, kind och mund Och rosor, löjen - allting der Så ljuft-vältaligt hvarje stund Om hjertats renhet vittne bär, Om frid med allt på jordensrund, Om kärlek, som blott oskuld är.

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1109-2238 (2015-01-17 17:34:42)

Tiu ĉi poemo troviĝas en http: // petercochran. files. wordpress. com/ 2009/ 03/ hebrew_ melodies. pdf .

Traduko de la Angla poemo “She Walks in Beauty” de Lord George Gordon Byron (∗1788-01-22 – †1824-04-18) en la Svedan de Theodor Lind.

Arg-1109-2254 (2015-01-17 17:43:44)

Tiu ĉi poem-traduko troviĝas en la retejo http: // runeberg. org/ hebrmel/ sub http: // runeberg. org/ hebrmel/ 0011. html ..

http://www.poezio.net/version?poem-id=1109&version-id=,2238,2254